# LimiX: Unleashing Structured-Data Modeling Capability for Generalist Intelligence

### v1

@@ Results.
As shown by critical difference diagrams, i.e. Figures 15, 16, 17 and 18, LimiX always achieves the best performance, outperforming strong baselines like AutoGluon, XGBoost, and TabPFN-v2. Quantitative results in Tables 13, 19, 15 and 17 also confirms that LimiX achieves state-of-the-art regression performance, leading in both normalized RMSE and R2.
@@
Similar to the subgroup analysis in classification, for regression, we also use the sample subgroups when building BCCO-REG to perform stratified analyses, whose criteria include the number of training samples, the ratio between the number of samples and features (length-to-width ratio), and the proportion of categorical features. Figure 20 shows that LimiX outperforms other methods across all subgroups compared with other methods. For Figures 20(b), 20(c) and 20(d), we can see that performance of Mitra substantially drops for the second subgroup, while LimiX do not exhibit a large drop.
@@
BCCO-REG
Mean Rank
LimiX 0.794±0.004 0.386±0.004 3.100 3.760
AutoGluon 0.781±0.001 0.398±0.000 3.700 3.660
TabPFN-v2 0.772±0.001 0.404±0.002 5.000 4.900
XGBoost 0.764±0.006 0.415±0.005 7.740 7.780
ET 0.757±0.005 0.431±0.006 9.960 9.900
RF 0.752±0.007 0.438±0.009 11.260 11.220
T2G-Former 0.743±0.011 0.442±0.012 12.860 12.820
ExcelFormer 0.743±0.015 0.443±0.012 11.460 11.520
CatBoost 0.741±0.026 0.427±0.020 10.520 10.360
DCN-v2 0.739±0.015 0.448±0.014 13.780 13.780
FT-Transformer 0.737±0.009 0.448±0.010 13.440 13.420
MLP-PLR 0.734±0.012 0.453±0.009 15.720 15.720
TabR 0.733±0.017 0.448±0.011 14.460 14.520
RealMLP 0.725±0.029 0.441±0.010 12.420 12.380
AutoInt 0.724±0.015 0.465±0.018 17.220 17.140
ResNet 0.720±0.009 0.468±0.010 15.020 14.980
TANGOS 0.719±0.010 0.468±0.009 15.680 15.680
LightGBM 0.715±0.019 0.423±0.005 7.980 7.940
MLP 0.701±0.013 0.487±0.010 16.940 17.000
SAINT 0.701±0.022 0.481±0.019 16.520 16.540
TabNet 0.670±0.024 0.516±0.024 20.640 20.580
Mitra 0.667±0.027 0.474±0.015 15.700 15.620
NODE 0.643±0.099 0.543±0.097 20.100 20.020
ModernNCA 0.598±0.195 0.471±0.038 15.180 15.080
SNN 0.434±0.111 0.720±0.072 25.200 25.200
GrowNet 0.201±0.182 0.864±0.110 26.840 26.820
DANets 0.005±0.008 0.979±0.004 28.440 28.480
SwitchTab 0.001±0.032 0.981±0.016 28.700 28.700
TabTransformer 0.000±0.016 0.981±0.008 28.600 28.640
DNNR -2.152±0.150 1.329±0.027 20.820 20.840
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 12 33 16 14 4 81
Missing Values (Ratio) 0 0 0 0 0 0 0.001
Categorical Features (Ratio) 0 0.16 0.65 0.242 0.264 0 0.967
Features w/ Missing Values (Ratio) 0 0 0 0.008 0.051 0 0.333
TALENT-REG
Mean Rank
LimiX 0.735±0.009 0.433±0.009 2.667 3.162
AutoGluon 0.722±0.006 0.448±0.009 4.313 4.414
XGBoost 0.710±0.006 0.462±0.005 6.768 6.697
LightGBM 0.707±0.004 0.461±0.004 7.485 7.455
CatBoost 0.700±0.015 0.471±0.018 9.667 9.606
RF 0.697±0.008 0.474±0.009 9.788 9.788
ET 0.696±0.008 0.476±0.007 9.242 9.192
TabPFN-v2 0.695±0.003 0.465±0.002 5.667 5.576
RealMLP 0.656±0.013 0.510±0.010 12.636 12.636
T2G-Former 0.656±0.019 0.512±0.017 13.364 13.303
MLP-PLR 0.653±0.017 0.521±0.017 15.242 15.202
ExcelFormer 0.653±0.016 0.512±0.012 12.444 12.485
TabR 0.651±0.023 0.516±0.021 14.434 14.394
FT-Transformer 0.648±0.019 0.519±0.018 15.000 14.949
AutoInt 0.636±0.020 0.538±0.021 18.152 18.162
ModernNCA 0.633±0.033 0.530±0.029 15.758 15.747
Mitra 0.602±0.053 0.547±0.022 14.909 14.899
TANGOS 0.592±0.027 0.547±0.013 16.202 16.172
TabNet 0.576±0.042 0.586±0.032 20.798 20.737
NODE 0.568±0.096 0.600±0.091 19.636 19.495
ResNet 0.562±0.095 0.550±0.023 15.697 15.667
MLP 0.556±0.025 0.564±0.008 16.677 16.717
SNN 0.344±0.138 0.777±0.069 24.485 24.505
DANets 0.005±0.009 0.998±0.004 27.475 27.495
TabTransformer 0.001±0.017 1.001±0.008 27.606 27.657
SwitchTab -0.002±0.027 1.002±0.013 27.980 28.010
GrowNet -0.182±0.319 0.920±0.120 26.162 26.192
DCN-v2 -0.361±0.161 0.818±0.182 14.717 14.747
SAINT -1.541±0.014 0.571±0.010 17.495 17.364
DNNR -9.172±5.278 2.528±0.814 22.525 22.566
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 11 79 30 49 4 266
Missing Values (Ratio) 0 0 0 0.007 0.034 0 0.283
Categorical Features (Ratio) 0 0.131 0.667 0.238 0.283 0 1
Features w/ Missing Values (Ratio) 0 0 0.013 0.043 0.161 0 0.875
CTR23
Mean Rank
LimiX 0.745±0.008 0.477±0.007 3.879 3.970
AutoGluon 0.725±0.001 0.497±0.001 4.818 4.727
TabPFN-v2 0.716±0.002 0.503±0.002 7.182 7.061
XGBoost 0.712±0.008 0.511±0.007 8.000 8.061
LightGBM 0.706±0.004 0.516±0.005 9.000 9.091
CatBoost 0.700±0.022 0.528±0.030 10.455 10.455
ET 0.697±0.007 0.535±0.008 9.758 9.818
RF 0.694±0.010 0.539±0.012 10.667 10.667
T2G-Former 0.674±0.013 0.544±0.014 13.061 13.030
MLP-PLR 0.672±0.013 0.553±0.014 15.000 15.030
TabR 0.671±0.014 0.543±0.014 14.182 14.121
DCN-v2 0.670±0.015 0.545±0.015 13.788 13.848
FT-Transformer 0.667±0.011 0.549±0.011 14.061 14.152
ModernNCA 0.667±0.027 0.550±0.027 15.121 14.788
ExcelFormer 0.665±0.014 0.556±0.013 11.727 11.818
RealMLP 0.661±0.013 0.549±0.011 13.788 13.818
AutoInt 0.655±0.015 0.568±0.018 16.970 16.970
SAINT 0.654±0.015 0.561±0.017 15.970 15.879
ResNet 0.645±0.018 0.587±0.017 13.667 13.727
TANGOS 0.642±0.014 0.586±0.013 15.818 15.848
Mitra 0.624±0.017 0.583±0.016 15.788 15.909
MLP 0.608±0.008 0.623±0.008 15.000 15.000
TabNet 0.605±0.046 0.623±0.041 19.606 19.667
NODE 0.568±0.105 0.666±0.096 19.636 19.485
SNN 0.369±0.073 0.834±0.048 24.727 24.727
GrowNet 0.185±0.184 0.944±0.113 26.606 26.576
DANets 0.001±0.010 1.052±0.005 28.000 28.000
TabTransformer 0.000±0.016 1.053±0.008 28.000 28.000
SwitchTab -0.006±0.025 1.057±0.013 29.061 29.061
DNNR -2.969±0.262 1.651±0.040 21.667 21.697
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 11 39 20 22 5 116
Missing Values (Ratio) 0 0 0 0.008 0.036 0 0.209
Categorical Features (Ratio) 0 0.143 0.842 0.312 0.345 0 1
Features w/ Missing Values (Ratio) 0 0 0 0.011 0.05 0 0.286
PFN-REG
Mean Rank
LimiX 0.692±0.012 0.461±0.011 4.143 4.964
AutoGluon 0.677±0.000 0.482±0.001 6.607 6.607
TabPFN-v2 0.676±0.004 0.475±0.002 7.250 7.036
XGBoost 0.671±0.007 0.487±0.007 8.679 8.679
RealMLP 0.664±0.012 0.484±0.010 9.429 9.357
LightGBM 0.664±0.008 0.496±0.008 9.714 9.643
CatBoost 0.661±0.014 0.498±0.014 10.607 10.571
ET 0.652±0.007 0.515±0.008 10.643 10.571
ExcelFormer 0.646±0.013 0.513±0.013 13.179 13.214
ModernNCA 0.645±0.021 0.508±0.015 14.286 14.464
RF 0.643±0.010 0.524±0.011 12.929 13.000
T2G-Former 0.640±0.027 0.503±0.022 13.143 13.036
MLP-PLR 0.638±0.017 0.513±0.019 14.429 14.500
DCN-v2 0.638±0.024 0.511±0.022 14.786 14.607
TabR 0.636±0.036 0.509±0.030 14.500 14.357
FT-Transformer 0.636±0.033 0.509±0.027 14.107 13.964
Mitra 0.630±0.014 0.531±0.016 12.821 12.857
AutoInt 0.610±0.031 0.539±0.033 18.214 18.107
ResNet 0.599±0.017 0.553±0.013 14.393 14.429
TANGOS 0.587±0.024 0.562±0.025 14.679 14.821
MLP 0.577±0.023 0.576±0.016 14.107 14.250
NODE 0.496±0.135 0.666±0.117 20.071 19.893
TabNet 0.430±0.216 0.643±0.106 21.107 21.036
SNN 0.378±0.079 0.756±0.056 24.179 24.214
GrowNet 0.096±0.141 0.942±0.080 25.679 25.643
DANets 0.001±0.012 0.997±0.006 26.893 26.929
TabTransformer -0.021±0.029 1.007±0.014 28.357 28.393
SwitchTab -0.026±0.034 1.011±0.017 27.429 27.464
DNNR -7.850±0.583 2.224±1.125 22.286 22.429
SAINT -8.088±0.020 0.612±0.014 16.357 15.964
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 16 176 66 93 3 376
Missing Values (Ratio) 0 0 0.089 0.029 0.076 0 0.335
Categorical Features (Ratio) 0 0.209 0.949 0.355 0.381 0 1
Features w/ Missing Values (Ratio) 0 0 0.211 0.059 0.169 0 0.841
@@ 7.3 Missing Value Imputation
The ubiquity of missing values in tabular data is harmful to downstream tasks or statistical analyses (Lin & Tsai, 2020). Meanwhile, the problem of missing value imputation poses a bigger challenge compared with standard classification and regression tasks since it requires modeling of the joint distribution of $P(X,Y)$ instead of $P(Y|X)$ only. A noteworthy capability of LimiX is missing value imputation, which naturally stems from the mask modeling during pretraining. Although there are existing deep learning approaches for missing value imputation (Jarrett et al., 2022; Zhang et al., 2025), they all require additional training on unseen datasets. In contrast, LimiX is the first model to be capable of missing value imputation on unseen datasets via in-context learning without additional training. This brings great convenience to the usage in downstream tasks.
We conduct experiments on the following datasets: Analcatdata BroadwayMult (Simonoff, 2003; Vanschoren, 2014a), Early Stage Diabetes 2020 (Islam et al., 2019; 2020), Forty Soybean Cultivars from Subsequent Harvests (de Oliveira et al., 2023; Rodrigues de Oliveira & Mario Zuffo, 2023), HCC survival (Santos et al., 2015b; a), Vehicle (Siebert, 1987; Mowforth & Shepherd, 1987), Eucalyptus (van Rijn, 2014), and Z-Alizadeh Sani (Alizadehsani et al., 2013). To simulate the scenario of missing value imputation, we randomly mask a fixed proportion $\alpha$ of cells in ${\mathbf{x}}_{te}$ and try to recover values of these cells. In our experiments, we set $\alpha=0.05$. Baselines include: Mean/Mode (use mean for continuous features and mode for categorical features), k-nearest neighbors (KNN), MICE (Van Buuren & Groothuis-Oudshoorn, 2011), MissForest (Stekhoven & Bühlmann, 2012), GAIN (Yoon et al., 2018), MIWAE (Mattei & Frellsen, 2019), and HyperImpute (Jarrett et al., 2022). For continuous features, we normalize each feature using MinMaxScaler and calculate RMSE between imputed values and ground truth values after normalization as the evaluation metric. For categorical features, the error rate is used as the evaluation metric. Table 21 shows that LimiX consistently outperforms previous baselines, all of which require additional training or fitting.

### v2

@@ Results.
As shown by critical difference diagrams, i.e. Figures 15, 16, 17, 18 and 19, LimiX-16M always achieves the best performance, outperforming strong baselines like AutoGluon, XGBoost, and TabPFN-v2. Quantitative results in Tables 13, 19, 15, 17 and 21 also confirms that LimiX-16M achieves state-of-the-art regression performance, leading in both normalized RMSE and R2. On the regression benchmarks, LimiX-2M exhibits predictive performance comparable to AutoGluon and TabPFN-v2, while requiring substantially less computational resources and runtime.
@@
Similar to the subgroup analysis in classification, for regression, we also use the sample subgroups when building BCCO-REG to perform stratified analyses, whose criteria include the number of training samples, the ratio between the number of samples and features (length-to-width ratio), and the proportion of categorical features. Figure 21 shows that LimiX-16M outperforms other methods across all subgroups compared with other methods. For Figures 21(b), 21(c) and 21(d), we can see that performance of Mitra substantially drops for the second subgroup, while LimiX-16M do not exhibit a large drop.
BCCO-REG
Mean Rank
LimiX-16M 0.794 0.386 3.340 4.120
AutoGluon 0.781 0.398 4.260 4.220
TabPFN-v2 0.772 0.404 5.480 5.380
LimiX-2M 0.785 0.392 5.720 5.600
XGBoost 0.764 0.415 8.520 8.560
LightGBM 0.715 0.423 8.760 8.720
ET 0.757 0.431 10.720 10.660
CatBoost 0.741 0.427 11.320 11.160
RF 0.752 0.438 12.080 12.040
ExcelFormer 0.743 0.443 12.260 12.320
RealMLP 0.725 0.441 13.240 13.200
T2G-Former 0.743 0.442 13.680 13.640
FT-Transformer 0.737 0.448 14.240 14.220
DCN-v2 0.739 0.448 14.600 14.600
TabR 0.733 0.448 15.300 15.360
ResNet 0.720 0.468 15.940 15.900
ModernNCA 0.598 0.471 16.060 15.960
TANGOS 0.719 0.468 16.620 16.620
Mitra 0.667 0.474 16.620 16.540
MLP-PLR 0.734 0.453 16.660 16.660
SAINT 0.701 0.481 17.440 17.460
MLP 0.701 0.487 17.860 17.920
AutoInt 0.724 0.465 18.120 18.040
NODE 0.643 0.543 21.000 20.920
TabNet 0.670 0.516 21.600 21.540
DNNR -2.152 1.329 21.780 21.800
SNN 0.434 0.720 26.200 26.200
GrowNet 0.201 0.864 27.840 27.820
DANets 0.005 0.979 29.440 29.480
TabTransformer -0.000 0.981 29.600 29.640
SwitchTab 0.001 0.981 29.700 29.700
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 12 33 16 14 4 81
Missing Values (Ratio) 0 0 0 0 0 0 0.001
Categorical Features (Ratio) 0 0.16 0.65 0.242 0.264 0 0.967
Features w/ Missing Values (Ratio) 0 0 0 0.008 0.051 0 0.333
TALENT-REG
Mean Rank
LimiX-16M 0.735 0.433 2.808 3.364
AutoGluon 0.722 0.448 4.818 4.929
TabPFN-v2 0.695 0.465 6.192 6.101
LimiX-2M 0.721 0.451 6.263 6.182
XGBoost 0.710 0.462 7.424 7.354
LightGBM 0.707 0.461 8.182 8.152
ET 0.696 0.476 9.970 9.919
CatBoost 0.700 0.471 10.424 10.354
RF 0.697 0.474 10.556 10.556
ExcelFormer 0.653 0.512 13.313 13.354
RealMLP 0.656 0.510 13.444 13.444
T2G-Former 0.656 0.512 14.202 14.141
TabR 0.651 0.516 15.283 15.242
DCN-v2 -0.361 0.818 15.576 15.616
FT-Transformer 0.648 0.519 15.848 15.798
Mitra 0.602 0.547 15.848 15.848
MLP-PLR 0.653 0.521 16.101 16.061
ResNet 0.562 0.550 16.616 16.586
ModernNCA 0.633 0.530 16.646 16.626
TANGOS 0.592 0.547 17.131 17.101
MLP 0.556 0.564 17.596 17.636
SAINT -1.541 0.571 18.384 18.253
AutoInt 0.636 0.538 19.040 19.051
NODE 0.568 0.600 20.515 20.374
TabNet 0.576 0.586 21.747 21.687
DNNR -9.172 2.528 23.505 23.545
SNN 0.344 0.777 25.465 25.485
GrowNet -0.182 0.920 27.131 27.162
DANets 0.005 0.998 28.434 28.455
TabTransformer 0.001 1.001 28.576 28.626
SwitchTab -0.002 1.002 28.949 28.990
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 11 79 30 49 4 266
Missing Values (Ratio) 0 0 0 0.007 0.034 0 0.283
Categorical Features (Ratio) 0 0.131 0.667 0.238 0.283 0 1
Features w/ Missing Values (Ratio) 0 0 0.013 0.043 0.161 0 0.875
CTR23
Mean Rank
LimiX-16M 0.745 0.477 4.061 4.152
AutoGluon 0.725 0.497 5.303 5.212
LimiX-2M 0.730 0.495 6.879 6.909
TabPFN-v2 0.716 0.503 7.636 7.515
XGBoost 0.712 0.511 8.697 8.758
LightGBM 0.706 0.516 9.727 9.818
ET 0.697 0.535 10.515 10.576
CatBoost 0.700 0.528 11.212 11.212
RF 0.694 0.539 11.485 11.455
ExcelFormer 0.665 0.556 12.515 12.606
T2G-Former 0.674 0.544 13.879 13.848
DCN-v2 0.670 0.545 14.515 14.576
ResNet 0.645 0.587 14.545 14.606
RealMLP 0.661 0.549 14.576 14.606
FT-Transformer 0.667 0.549 14.848 14.939
TabR 0.671 0.543 14.939 14.879
MLP-PLR 0.672 0.553 15.788 15.818
MLP 0.608 0.623 15.848 15.848
ModernNCA 0.667 0.550 15.909 15.576
Mitra 0.624 0.583 16.727 16.848
TANGOS 0.642 0.586 16.727 16.758
SAINT 0.654 0.561 16.758 16.667
AutoInt 0.655 0.568 17.848 17.818
NODE 0.568 0.666 20.545 20.394
TabNet 0.605 0.623 20.576 20.636
DNNR -2.969 1.651 22.636 22.667
SNN 0.369 0.834 25.667 25.697
GrowNet 0.185 0.944 27.606 27.576
DANets 0.001 1.052 28.970 28.970
TabTransformer 0.000 1.053 29.000 29.000
SwitchTab -0.006 1.057 30.061 30.061
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 11 39 20 22 5 116
Missing Values (Ratio) 0 0 0 0.008 0.036 0 0.209
Categorical Features (Ratio) 0 0.143 0.842 0.312 0.345 0 1
Features w/ Missing Values (Ratio) 0 0 0 0.011 0.05 0 0.286
PFN-REG
Mean Rank
LimiX-16M 0.692 0.461 4.607 5.464
LimiX-2M 0.696 0.458 5.250 5.250
AutoGluon 0.677 0.482 7.321 7.321
TabPFN-v2 0.676 0.475 7.857 7.607
XGBoost 0.671 0.487 9.464 9.464
RealMLP 0.664 0.484 10.250 10.179
LightGBM 0.664 0.496 10.464 10.393
CatBoost 0.661 0.498 11.429 11.393
ET 0.652 0.515 11.500 11.429
Mitra 0.630 0.531 13.679 13.714
RF 0.643 0.524 13.821 13.893
T2G-Former 0.640 0.503 13.964 13.857
ExcelFormer 0.646 0.513 14.036 14.071
FT-Transformer 0.636 0.509 14.964 14.821
MLP 0.577 0.576 14.964 15.107
ModernNCA 0.645 0.508 15.107 15.286
ResNet 0.599 0.553 15.286 15.321
MLP-PLR 0.638 0.513 15.321 15.393
TabR 0.636 0.509 15.393 15.250
TANGOS 0.587 0.562 15.607 15.750
DCN-v2 0.638 0.511 15.607 15.429
SAINT -8.088 0.612 17.214 16.821
AutoInt 0.610 0.539 19.143 19.036
NODE 0.496 0.666 20.964 20.786
TabNet 0.430 0.643 22.071 22.000
DNNR -7.850 2.224 23.286 23.429
SNN 0.378 0.756 25.179 25.214
GrowNet 0.096 0.942 26.607 26.571
DANets 0.001 0.997 27.893 27.929
SwitchTab -0.026 1.011 28.393 28.429
TabTransformer -0.021 1.007 29.357 29.393
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 16 176 66 93 3 376
Missing Values (Ratio) 0 0 0.089 0.029 0.076 0 0.335
Categorical Features (Ratio) 0 0.209 0.949 0.355 0.381 0 1
Features w/ Missing Values (Ratio) 0 0 0.211 0.059 0.169 0 0.841
Table 21: Regression results on the TabArena-REG benchmark. The best scores are shown in bold.
TabArena-REG
Mean Rank
LimiX-16M 0.796 0.406 2.923 2.923
AutoGluon 0.791 0.414 3.154 3.154
LimiX-2M 0.788 0.413 3.846 3.846
TabPFN-v2 0.777 0.422 5.385 5.385
CatBoost 0.774 0.431 6.231 6.231
XGBoost 0.778 0.430 6.231 6.231
LightGBM 0.771 0.435 6.769 6.769
RF 0.758 0.456 9.769 9.769
ET 0.746 0.464 10.231 10.154
TabR 0.729 0.476 13.692 13.692
ExcelFormer 0.681 0.521 13.923 13.923
NODE 0.665 0.542 14.538 14.538
TANGOS 0.673 0.534 15.077 15.077
DCN-v2 0.718 0.486 15.385 15.462
RealMLP 0.719 0.487 15.385 15.385
Mitra 0.666 0.538 16.615 16.615
ModernNCA 0.712 0.496 17.231 17.231
MLP-PLR 0.714 0.496 17.231 17.231
AutoInt 0.705 0.503 17.923 17.923
ResNet 0.687 0.521 18.538 18.538
MLP 0.694 0.516 18.538 18.538
SAINT 0.712 0.498 19.000 19.000
T2G-Former 0.677 0.526 19.385 19.385
TabNet 0.641 0.564 19.615 19.615
FT-Transformer 0.626 0.572 22.077 22.077
DNNR -0.368 1.058 24.538 24.538
SNN 0.451 0.729 25.692 25.692
GrowNet 0.316 0.814 27.385 27.385
TabTransformer 0.016 0.993 29.333 29.333
DANets 0.012 0.998 29.462 29.462
SwitchTab 0.004 1.002 29.923 29.923
Table 22: Statistical profile of the benchmark TabArena-REG, where Q10, Q50, and Q90 correspond to the 10%, 50%, and 90% quantiles, respectively; for categorical feature statistics, we only consider features that are either string-typed or have fewer than 10 unique values.
Statistics
Metric Q10 Q50 Q90 Mean Std Min Max
# Features 6 9 67.8 92.23 280.67 5 1024
Missing Values (Ratio) 0 0 0 0 0 0 0
Categorical Features (Ratio) 0 0.333 0.619 0.287 0.299 0 1
Features w/ Missing Values (Ratio) 0 0 0 0 0 0 0
@@
(a) R2 on the TabArena-REG benchmark.
(b) RMSE on the TabArena-REG benchmark.
Figure 19: Critical difference diagrams on TabArena-REG benchmark.
@@ 7.3 Missing Value Imputation
The ubiquity of missing values in tabular data is harmful to downstream tasks or statistical analyses (lin2020missing). Meanwhile, the problem of missing value imputation poses a bigger challenge compared with standard classification and regression tasks since it requires modeling of the joint distribution of $P(X,Y)$ instead of $P(Y|X)$ only. A noteworthy capability of LimiX models is missing value imputation, which naturally stems from the mask modeling during pretraining. Although there are existing deep learning approaches for missing value imputation (jarrett2022hyperimpute; zhang2025diffputer), they all require additional training on unseen datasets. In contrast, LimiX models are the first to perform missing value imputation on unseen datasets via in-context learning without additional training. This brings great convenience to the usage in downstream tasks.
We conduct experiments on the following datasets: Analcatdata BroadwayMult (simonoff2003analyzing; openml_analcatdata), Early Stage Diabetes 2020 (islam2019likelihood; early_stage_diabetes_risk_prediction_529), Forty Soybean Cultivars from Subsequent Harvests (de2023dataset; forty_soybean_cultivars_from_subsequent_harvests_913), HCC survival (santos2015new; hcc_survival_423), Vehicle (siebert1987vehicle; statlog_(vehicle_silhouettes)_149), Eucalyptus (openml_eucalyptus), and Z-Alizadeh Sani (z-alizadeh_sani_412). To simulate the scenario of missing value imputation, we randomly mask a fixed proportion $\alpha$ of cells in ${\mathbf{x}}_{te}$ and try to recover values of these cells. In our experiments, we set $\alpha=0.05$. Baselines include: Mean/Mode (use mean for continuous features and mode for categorical features), k-nearest neighbors (KNN), MICE (van2011mice), MissForest (stekhoven2012missforest), GAIN (yoon2018gain), MIWAE (mattei2019miwae), and HyperImpute (jarrett2022hyperimpute). For continuous features, we normalize each feature using MinMaxScaler and calculate RMSE between imputed values and ground truth values after normalization as the evaluation metric. For categorical features, the error rate is used as the evaluation metric. Table 23 shows that LimiX-16M consistently outperforms previous baselines, all of which require additional training or fitting.